OpenStudy (anonymous):

The following sequence is geometric. What is the next term? 12, 36, 108, 324 ... 412 608 500 972

Parth (parthkohli):

You are again and again multiplying 3. 324 * 3

Parth (parthkohli):

Mayhaps cwrw is better at explaining :S

OpenStudy (cwrw238):

common ratio = 36/12 = 3 each term is 3* previous so next term is what parthkohli said

OpenStudy (cwrw238):

you explained it well parthkohli

OpenStudy (anonymous):

just multiply it by 3...

OpenStudy (anonymous):

Find the common ratio first: 32/12 = 3 Now the formula to be used is: \[Term = a.r^{n-1}\] Which term you are finding?? It is 5th So, n= 5 So, Term = 12(3)^{5-1} = 12(3)^4 = 12*81 Multiply to get the answer..

OpenStudy (anonymous):

The common ratio between the numbers is 3, because 36 divided by 12 is 3, and 108 divided by 36 is 3. It continues this way. So all you have to do to find the next number in the series is multiply the last number by 3. 324 x 3 = 972 The answer is: D. 972